# Revisiting Multi-User Downlink in IEEE 802.11ax: A Designers Guide to MU-MIMO

### v1

@@ I Introduction
IEEE 802.11ax (Wi-Fi 6) marked a significant evolution milestone via the introduction of Multi-User (MU) communication modes (in contrast with legacy Single-User (SU) communication) for both Uplink (UL) and Downlink (DL) in tri-band (2.4/5/6 GHz) [1]. For the Uplink, this implies the use of trigger-based OFDMA; in this article, we focus solely on DL Multi-User (MU) Multiple Input Multiple Output (MU-MIMO). Legacy Single-User MIMO (SU-MIMO) - the precursor to MU-MIMO - laid the groundwork by allowing transmission of multiple spatial streams from an access point (AP) equipped with multiple antennas to a single client device on downlink. With the proliferation of wireless client devices, a single Wi-Fi network access point (AP) can have many (typically $10-20$) associated stations (STAs) [2, 3]. With multi-antenna clients  [^1] However, the number of antennas at the AP always exceeds the number of antennas at a client., it is feasible via DL Transmit Beamforming (TxBF) at the AP to send multiple streams to multiple STAs simultaneously (DL MU-MIMO).
A typical configuration [4, 5] such as Fig. 1 assumes an 8 x 8 AP (e.g., NetGear RAXE500) and 2 x 2 STAs (e.g., iPhone 15 and MacBook Air), implying that a single downlink transmission opportunity can potentially send a total of $8$ spatial streams  [^2] Note that Wi-Fi 5 (IEEE 802.11ac) included support for MU-MIMO but limited to 4 streams on only 5 GHz downlink operation; whereas Wi-Fi 6 supports up to 8-stream on 2.4/5/6 GHz uplink/downlink operations. to a selected sub-set of clients (e.g. 2 streams each to 4 STAs). While DL SU-MIMO results in scaling of per-user throughput as a result of multi-stream transmission, its benefits are limited by the fact that most clients support either 1 or 2 spatial streams (i.e., a total of 2-stream transmissions in DL SU-MIMO in Fig. 1). By contrast, it is evident that in dense overlapped network scenarios - such as the enterprise or residential cluster - DL MU-MIMO provides a natural pathway to increasing network efficiency (aggregate network throughput) by enabling simultaneous transmissions of multiple streams to multiple clients (i.e., a total of 8-stream transmissions in DL MU-MIMO in Fig. 1), with appropriate choice of the user sub-set and TxBF to minimize inter-user/inter-stream interference.
@@ II-A DL SU/MU-MIMO channel sounding
DL SU-MIMO is indicated by feedback type 0 in the High-Efficiency (HE) MIMO control field. As Fig. 2 shows, its channel sounding process consists of four major steps:
@@
By contrast, DL MU-MIMO, indicated by feedback type 1 in the HE MIMO control field, follows the similar channel sounding protocols as the SU-MIMO, however, several major differences exist:
@@ II-B CSI Overhead Comparison
CSI overhead in DL SU/MU-MIMO can be calculated based on the CSI frame format indicating each sub-field size, as shown in Fig. 2. In particular, CSI overhead is dominated by HE compressed BF feedback that contains the HE compressed BF report (as well as the extra sub-field - HE MU Exclusive BF report in MU-MIMO). The compressed BF report contains the compressed CSI for each sub-carrier, i.e., the V-matrix or steering/precoding matrix [^5] The Null-steering step based on zero-forcing (ZF) and minimum mean square error (MMSE) approaches [12, 14], used for precoding in DL MU-MIMO are not implemented in real AP products [10, 7] because those can be implemented only if the full CSI is obtained, whereas the fedback V-matrix provides only partial CSI. Besides, null-steering step incurs additional computational complexity and thus chipset cost for the AP. used for digital beamforming. V-matrix is obtained by a) applying the singular value decomposition (SVD) to the full CSI, and b) compressing it to specific Givens rotation angles to reduce the amount of required bits. The compressed size of the V-matrix depends on the total number of Givens rotation angles as well as the number of bits used to quantize each angle, as defined in IEEE 802.11ax specification. In general, the larger the V-matrix dimension, the more the number of angles. Meanwhile, the number of bits to quantize each angle is indicated by the sub-field of codebook information choice with 1 bit in the HE MIMO Control field. Thus both SU- and MU- have two codebook information choices [1], however, MU-MIMO uses more bits than SU-MIMO to quantize a single angle by using the same codebook information. For instance, if codebook information bit is set to 0, the number of bits to quantize an angle in SU-MIMO is 4 or 2 while they are 7 or 5 in MU-MIMO [1], implying that the compressed V-matrix in MU-MIMO has larger overhead compared to SU-MIMO. In addition, the HE compressed BF report size also scales with the number of spatial streams and the number of sub-carriers. The MU-Exclusive BF report in MU-MIMO contains the delta SNR per sub-carrier, which represents the difference from the average SNR. The MU-Exclusive BF report represents the spatial characteristics for each sub-carrier caused by the environment, the size of which scales with the number of subcarriers. Since the 802.11ax specification does not detail how this information is exploited in the design of the beamformer, its implementation is chip vendor dependent.
@@
Assuming the 8 x 8 AP, 1 x 1 STA and 20 MHz (242 sub-carriers) bandwidth as in Fig. 3, the effective channel capacity does not grow linearly with the number of STAs. In particular, the effective channel capacity under codebook info 1 is greatly reduced when the number of STAs reaches 8. This can be explained by following reasons:
@@ III-A Clustered-based multi-path channel model
We use the class of cluster-based multipath fading channels to model the practical environmental conditions for indoor Wi-Fi downlink operation. Such models were introduced by Saleh and Valenzuela, and extended/elaborated upon by many other researchers [14]. In particular, IEEE 802.11ax indoor channel model [11] is a typical cluster-based channel model that we have adapted by incorporating a parameter for user separation, as shown in Fig. 4. IEEE 802.11ax indoor channel model represents the propagation environment as a collection of scatterers grouped into clusters, where each cluster represents objects in the vicinity that act as a forward scattering source of rays that reach the receiver. Such clusters are typically represented via spatio-temporal models that capture the spatial characteristics of the environment, such as the transmit/receive antenna correlation and the distribution of objects, etc.
Figure 4: DL SU (STA 1) and MU (STA 1 + 2) in Line-of-sight Scenario under Adapted IEEE 802.11ax Indoor Channel Model.
A particular impact on our results arises from distinction between Line-of-sight (LoS) and Non-line-of-sight (NLoS) scenarios as defined by 11ax channel model specification, depending on the relationship between the breakpoint distance  [^8] The distance that separates LoS and NLoS scenarios with different path loss exponents. and the AP-STA separations[11]:
@@
Figure 5: Channel Capacity impacted by Spatial Correlation. 8 x 8 AP (ULA antenna spacing is half wavelength), 1 x 1 STAs, 20 dBm Transmit Power, 20 MHz Bandwidth, -174 dBm Noise Power Spectrum Density.
@@ III-B Spatial Correlation
Fig. 4 shows a two-cluster scenario including an 8 x 8 uniform linear array (ULA)-based AP and two 1 x 1 STAs. AP transmits to STA 1 if MU-MIMO is turned off and to both STA 1 and STA 2 if MU-MIMO is turned on. The spatial geometry of STAs is characterized by their angle of departure (AoD), i.e., LoS AoD to STA 1 and LoS AoD to STA 2. The user separation between STA 1 and 2 is defined as the difference between LoS AoD to STA 2 and STA 1, respectively. To investigate the impact of user separation, we fix the angular geometry of cluster 1999Cluster 1’s NLoS mean AoD equals the LoS AoD to STA 1; Cluster 2’s NLoS mean AoD equals the LoS AoD to STA 2. and STA 1, i.e., LoS AoD to STA 1 is set to $0^{o}$, and the LoS AoD to STA 2 varied between $0^{o}$ and $180^{o}$, thus the user separation between STA 1 and 2 ranges between $0^{o}$ and $180^{o}$.
@@ III-B1 Dominant feature in the LoS scenario - User separation
As the LoS distance is typically small due to the small breakpoint distance, e.g., within 10 meters, spatial correlation in the LoS scenario is not sensitive to the LoS distance variation. Thus user separation is the single dominant feature that we explore in the LoS scenario, which is shown in Fig. 5. Consider a set of LoS scenarios where the distance is 8 meters and the granularity of user separation is $2^{o}$, resulting in a total of 90 user separation scenarios. DL SU channel capacity dominates MU in 14% scenarios of which 12% scenarios lie in $0-12^{o}$ and $168-180^{o}$ user separation regions. Note that DL MU-MIMO channel capacity over $0-180^{o}$ user separation exhibits a symmetric channel capacity pattern, that can be attributed to the ULA characteristics where the LoS transmit/receive steering vectors of two STAs are identical at $0^{o}$ or $180^{o}$ user separation. Then, their dominant LoS CSI component determined by LoS transmit/receive steering vectors are also close for user separation regions close to $0^{o}$ or $180^{o}$. As a result, the corresponding V-matrices become highly correlated, incurring significantly higher inter-user interference than other user separation regions.
Figure 6: Guideline Table for DL MU-MIMO with 8 x 8 AP under Adapted IEEE 802.11ax Channel Model.
@@ IV Design Guideline for DL MU-MIMO
We now provide some practical design guidelines for DL MU-MIMO operation with 8 x 8 AP based on the effective channel capacity that unifies the underlying factors discussed in Section II and III. For the same setup as Fig. 5 used to obtain the channel capacity is now modified to derive the effective channel capacity as in Fig. 3. All results were implemented in Matlab using indoor MIMO WLAN channel models created by Schumacher et al, [15]. Due to page limits, we only provide guidelines for the 2-user case; this will be extended to cover other typical operational scenarios and reported elsewhere. Investigating more users results in more complicated scenarios to be categorized (e.g., a 3-user case will be characterized by 3 different user separations between each pair of users) and is left for future research work.
As the main features regarding CSI overhead are codebook information for BF compression (i.e., codebook info 0 and 1) and STA MIMO dimensions (i.e., 1 x 1 and 2 x 2 STA), there are a total of 4 different operational scenarios regarding CSI overhead. Meanwhile, spatial correlation is characterized by 2 typical user separation regions for LoS scenario, while it is characterized by 2 typical distance regions in the NLoS scenario. Thus there are also a total of 4 different operational scenarios regarding spatial correlation. As a result, we provide guidelines for 16 scenarios unifying both CSI overhead and spatial correlation, as shown in Fig. 6. Our conclusion for the 2-user case is that among these 16 scenarios, DL MU-MIMO can be turned on in 7 (43.75%). According to the guideline table, DL MU-MIMO can be turned on in the following scenarios:
@@
  • Sufficiently large distance in NLoS.
Otherwise, DL MU-MIMO is suggested to be turned off, i.e., switch to DL SU-MIMO. Note that the condition for turning on DL MU-MIMO is more stringent for the 2 x 2 STA case, compared to the 1 x 1 STA case. This is because each spatial stream in the 2 x 2 STA case suffers more from interfering streams (self-interference from another stream for same STA and/or streams from another STA) than the 1 x 1 STA case (only one interfering streams from another STA). Thus compared to the 1 x 1 STA case, MU-MIMO effective channel capacity is less likely to exceed SU-MIMO in the 2 x 2 STA case.

### v2

@@ I Introduction
IEEE 802.11ax (Wi-Fi 6) marked a significant evolution milestone via the introduction of Multi-User (MU) communication modes (in contrast with legacy Single-User (SU) communication) for both Uplink (UL) and Downlink (DL) in tri-band (2.4/5/6 GHz) [1]. For the Uplink, this implies the use of trigger-based OFDMA; in this article, we focus solely on DL Multi-User (MU) Multiple Input Multiple Output (MU-MIMO). Legacy Single-User MIMO (SU-MIMO) - the precursor to MU-MIMO - laid the groundwork by allowing transmission of multiple spatial streams from an access point (AP) equipped with multiple antennas to a single client device on downlink. With the proliferation of wireless client devices, a single Wi-Fi network access point (AP) can have multiple associated stations (STAs) [2, 3]. With multi-antenna clients  [^1] However, the number of antennas at the AP always exceeds the number of antennas at a client., it is feasible via DL Transmit Beamforming (TxBF) at the AP to send multiple streams to multiple STAs simultaneously (DL MU-MIMO).
A typical configuration [4, 5] such as Fig. 1 assumes an 8 x 8 AP (e.g., NetGear RAXE500) and 2 x 2 STAs (e.g., iPhone 15 and MacBook Air), implying that a single downlink transmission opportunity can potentially send a total of $8$ spatial streams  [^2] Note that Wi-Fi 5 (IEEE 802.11ac) included support for MU-MIMO but limited to 4 streams on only 5 GHz downlink operation; whereas Wi-Fi 6 supports up to 8-stream on 2.4/5/6 GHz uplink/downlink operations. to a selected sub-set of clients, e.g. each 2 streams to each selected four STAs. While DL SU-MIMO results in scaling of per-user throughput as a result of multi-stream transmission, its benefits are limited by the fact that most clients support either 1 or 2 spatial streams (i.e., a total of 2-stream transmissions in DL SU-MIMO in Fig. 1). By contrast, it is evident that in dense overlapped network scenarios - such as the enterprise or residential cluster - DL MU-MIMO provides a natural pathway to increasing network efficiency (aggregate network throughput) by enabling simultaneous transmissions of multiple streams to multiple clients (i.e., a total of 8-stream transmissions in DL MU-MIMO in Fig. 1), with appropriate choice of the user sub-set and TxBF to minimize inter-user/inter-stream interference.
@@ II-A DL SU/MU-MIMO channel sounding
DL SU-MIMO is indicated by codebook info 0 in the High-Efficiency (HE) MIMO control field. As Fig. 2 shows, its channel sounding process consists of four major steps:
@@
By contrast, DL MU-MIMO, indicated by codebook info 1 in the HE MIMO control field, follows the similar channel sounding protocols as the SU-MIMO, however, several major differences exist:
@@ II-B CSI Overhead Comparison
CSI overhead in DL SU/MU-MIMO can be calculated based on the CSI frame format indicating each sub-field size, as shown in Fig. 2. In particular, CSI overhead is dominated by HE compressed BF feedback that contains the HE compressed BF report (as well as the extra sub-field - HE MU Exclusive BF report in MU-MIMO). The compressed BF report contains the compressed CSI for each sub-carrier, i.e., the V-matrix or steering/precoding matrix [^5] The Null-steering step based on zero-forcing (ZF) and minimum mean square error (MMSE) approaches [12, 14], used for precoding in DL MU-MIMO are not implemented in real AP products [10, 7] because those can be implemented only if the full CSI is obtained, whereas the feedback V-matrix provides only partial CSI. Besides, null-steering step incurs additional computational complexity and thus chipset cost for the AP. used for digital beamforming. V-matrix is obtained by a) applying the singular value decomposition (SVD) to the full CSI, and b) compressing it to specific Givens rotation angles to reduce the amount of required bits. The compressed size of the V-matrix depends on the total number of Givens rotation angles as well as the number of bits used to quantize each angle, as defined in IEEE 802.11ax specification. In general, the larger the V-matrix dimension, the more the number of angles. Meanwhile, the number of bits to quantize each angle is indicated by the sub-field of codebook information choice with 1 bit in the HE MIMO Control field. Thus both SU- and MU- have two codebook information choices [1], however, MU-MIMO uses more bits than SU-MIMO to quantize a single angle by using the same codebook information. For instance, if codebook information bit is set to 0, the number of bits to quantize an angle in SU-MIMO is 4 or 2 while they are 7 or 5 in MU-MIMO [1], implying that the compressed V-matrix in MU-MIMO has larger overhead compared to SU-MIMO. In addition, the HE compressed BF report size also scales with the number of spatial streams and the number of sub-carriers. The MU-Exclusive BF report in MU-MIMO contains the delta SNR per sub-carrier, which represents the difference from the average SNR. The MU-Exclusive BF report represents the spatial characteristics for each sub-carrier caused by the environment, the size of which scales with the number of subcarriers. Since the 802.11ax specification does not detail how this information is exploited in the design of the beamformer, its implementation is chip vendor dependent.
@@
Assuming an 8 x 8 AP, 1 x 1 STA(s), and 20 MHz channel bandwidth as in Fig. 3, the effective channel capacity does not grow linearly with the number of STAs. In particular, the effective channel capacity under codebook info 1 is greatly reduced when the number of STAs reaches 8. This can be explained by following reasons:
@@ III-A Clustered-based multi-path channel model
We use the class of cluster-based multipath fading channels to model the practical environmental conditions for indoor Wi-Fi downlink operation. Such models were introduced by Saleh and Valenzuela, and extended/elaborated upon by many other researchers [14]. In particular, IEEE 802.11ax indoor channel model [11] is a typical cluster-based channel model that we have adapted by incorporating a parameter for user separation, as shown in Fig. 4. IEEE 802.11ax indoor channel model represents the propagation environment as a collection of scatterers grouped into clusters, where each cluster represents objects in the vicinity that act as a forward scattering source of rays that reach the receiver. Such clusters are typically represented via spatial-temporal models that capture the spatial characteristics of the environment, such as the transmit/receive antenna correlation and the distribution of objects, etc.
Figure 4: Modified IEEE 802.11ax Indoor Channel Model: DL SU (STA 1) and MU (STA 1 + 2) in Line-of-sight Scenario.
A particular impact on our results arises from distinction between Line-of-sight (LoS) and Non-line-of-sight (NLoS) scenarios as defined by 11ax channel model specification, depending on the relationship between the breakpoint distance  [^8] The breakpoint distance is defined as the distance that separates LoS and NLoS scenarios by characterizing different path loss exponents. and the distance between AP and STA(s) [11]:
@@
Figure 5: Channel Capacity impacted by Spatial Correlation. 20 dBm Transmit Power, 20 MHz Bandwidth, -174 dBm/Hz Noise Power Spectrum Density.
@@ III-B Spatial Correlation
Fig. 4 includes an 8 x 8 uniform linear array (ULA)-based AP whose ULA antenna spacing is half wavelength as well as two 1 x 1 STAs. The sake of using a 2-user example here is to provide key insights for readers. Extending to a larger user number will be discussed in the next section. Thus AP transmits to STA 1 if MU-MIMO is turned off and to both STA 1 and STA 2 if MU-MIMO is turned on. The spatial geometry of STAs is characterized by their angle of departure (AoD), i.e., LoS AoD to STA 1 and LoS AoD to STA 2. The user separation between STA 1 and 2 is defined as the difference between LoS AoD to STA 2 and STA 1, respectively. To investigate the impact of user separation, we fix the angular geometry of cluster 1999Cluster 1’s NLoS mean AoD equals the LoS AoD to STA 1; Cluster 2’s NLoS mean AoD equals the LoS AoD to STA 2. and STA 1, i.e., LoS AoD to STA 1 is set to $0^{o}$, and the LoS AoD to STA 2 varied between $0^{o}$ and $180^{o}$, thus the user separation between STA 1 and 2 ranges between $0^{o}$ and $180^{o}$.
@@ III-B1 Dominant feature in the LoS scenario - User separation
Due to the small breakpoint distance, e.g., 10 meters, spatial correlation in the LoS scenario is not sensitive to the distance variation. Thus user separation is the single dominant feature that we explore in the LoS scenario, which is shown in Fig. 5. Consider a set of LoS scenarios where the distance is 8 meters and the granularity of user separation is $2^{o}$, resulting in a total of 90 user separation scenarios. DL SU channel capacity dominates MU in 14% scenarios of which 12% scenarios lie in $0-12^{o}$ and $168-180^{o}$ user separation regions. Note that DL MU-MIMO channel capacity over $0-180^{o}$ user separation exhibits a symmetric channel capacity pattern, that can be attributed to the ULA characteristics where the LoS transmit/receive steering vectors of two STAs are identical at $0^{o}$ or $180^{o}$ user separation. Then, their dominant LoS CSI component determined by LoS transmit/receive steering vectors are also close for user separation regions close to $0^{o}$ or $180^{o}$. As a result, the corresponding V-matrices become highly correlated, incurring significantly higher inter-user interference than other user separation regions.
Figure 6: A 4-user Guideline Table for 8 x 8 AP under Modified IEEE 802.11ax Channel Model.
@@ IV Design Guideline for DL MU-MIMO
This section provides practical design guidelines that unify the underlying factors discussed in Section II and III. For the same setup as Fig. 5 used to obtain the channel capacity is now modified to derive the effective channel capacity as in Fig. 3. Meanwhile, We extend to a 4-user MU-MIMO operation, i,e, the user sub-set selection size is 4, indicating that upto 4 out of $X\geq 4$ STAs are selected if MU-MIMO is turned on. As the 4-user spatial correlation (where each is characterized by LoS/NLoS, user separation, and distance) results in a large set of scenario combinations, we thereby provide some typical scenarios due to the page limit. It should also be noted that real indoor channels might differ from the used channel model, that is, the exact spatial correlation threshold, such as $12^{o}$ user separation and $38$ meter distance in Section III, used for turning on/off MU-MIMO might be different. However, real channels should have the same guideline trend as the used channel model under each operational scenario (without specifying specific thresholds) defined in Fig. 6. All results were implemented in Matlab using indoor MIMO WLAN channel models created by Schumacher et al, [15].
As the main features regarding CSI overhead are codebook information for BF compression (i.e., codebook info 0 and 1) and STA MIMO dimensions (i.e., 1 x 1 and 2 x 2 STA), there are a total of 4 operational scenarios regarding CSI overhead. Meanwhile, we provide 5 typical operational scenarios (i.e., 2 LoS and 3 NLoS scenarios  [^10] For the operational scenario of two at small distances and two at large distances, AP is assumed to serve one of the STAs at small distances if MU-MIMO is turned off.) regarding spatial correlation. As a result, we provide guidelines for 20 scenarios unifying both CSI overhead and spatial correlation, as shown in Fig. 6. Our conclusion for the 2-user case is that among these 20 scenarios, DL MU-MIMO can be turned on in 9 (45%). According to the guideline table, DL MU-MIMO can be turned on in the following scenarios:
@@
  • 1 x 1 STAs in NLoS;
  • 2 x 2 STAs with codebook info 0 and large distances in NLoS.
Otherwise, DL MU-MIMO is suggested to be turned off, i.e., switch to DL SU-MIMO. Note that the condition for turning on DL MU-MIMO is more stringent for the 2 x 2 STA case, compared to the 1 x 1 STA case. This is because each spatial stream in the 2 x 2 STA case suffers more from interfering streams (self-interference from another stream for the same STA and/or streams from another STA) than the 1 x 1 STA case (only one interfering stream from another STA). Thus compared to the 1 x 1 STA case, MU-MIMO effective channel capacity is less likely to exceed SU-MIMO in the 2 x 2 STA case.